OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − 3y = 3 5x − 4y = 4 −1 0 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system

OpenStudy (anonymous):

think again with me!!

OpenStudy (anonymous):

you need to eliminate either x or y

OpenStudy (anonymous):

this is so hard i am so not good at this type of math ... it looks like they should both be canceled out

OpenStudy (anonymous):

ok! multiply the first equation by -5 and the second one by 2 what do u get?

OpenStudy (anonymous):

doing the math

OpenStudy (anonymous):

add the two after you multiply

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

no wait -1? -7 divided by 7 is negative one?

OpenStudy (anonymous):

good

OpenStudy (anonymous):

so i was correct?

OpenStudy (anonymous):

with -1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

find x now

OpenStudy (anonymous):

i thought that was what i was looking for

OpenStudy (anonymous):

y=-1 x=0

OpenStudy (anonymous):

here we need to find x not y! so the correct answer is 0

OpenStudy (anonymous):

to find x directly multiply equation 1 by -4 and equation 2 by 3 then add them

OpenStudy (campbell_st):

given your equations 2x - 3y = 3 (1) 5x - 4y = 4 (2) multiply equation (1) by 4 and equation 2 by (-3) and add the equations to eliminate y 8x - 12y = 12 -15x + 12y = -12 ----------------- -7x = 0 therefore x = 0